OpenStudy (josuelash4):

If the coordinates of point A are (8, 0) and the coordinates of point B are (3, 7), the y-intercept of AB is­­­_____ . If the coordinates of point D are (5, 5), the equation of line CD is y =____ x +____ .

OpenStudy (aakashtomar):

First, take out the equation of the line through two point form. \[y-y' = [(y' - y'')\div( x'- x'') ] \times ( x - x' ) \] where y' and y'' and x' and x'' are the y and x coordinates respectively of the points through which line passes. The slope intercept form of equation of line is \[y = mx + c\] where m is the slope and c is the y intercept. Try converting the equation you get into this form. \[( y' - y'') \div ( x' -x'') = m\] might help too.